OpenStudy (anonymous):

what is the simplified form of : 4x+2/x+5 - 3x+1/x+5 ?

OpenStudy (anonymous):

\[\frac{ 4x+2 }{ x+5 } - \frac{ 3x+1 }{ x+5}\]

OpenStudy (ivettef365):

that is the same as 4x + 2 - 3x - 1 ---------------- x+5

OpenStudy (ivettef365):

now combine like terms on the numerator

OpenStudy (anonymous):

i still dont get it

OpenStudy (ivettef365):

you can combine 4x - 3x = ??? and 2 - 1 = ???

OpenStudy (anonymous):

i dont get it

OpenStudy (ankit042):

@wolf_queen can you solve 2/5-3/5 ?

OpenStudy (anonymous):

i got it now

OpenStudy (ankit042):

K great!